OpenStudy (anonymous):

fan medal instagram For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8.

Vocaloid (vocaloid):

average rate of change: [f(8)-f(3)]/(8-3) = ?

OpenStudy (anonymous):

I don't get this concept sry @Vocaloid

Vocaloid (vocaloid):

f(x) = 0.01(2)^x f(8) = ?

OpenStudy (anonymous):

Also here r the answers 0.08 0.496 2.48 5

OpenStudy (anonymous):

irdk

Vocaloid (vocaloid):

f(x) = 0.01(2)^x let x = 8 what is 0.01(2)^x?

OpenStudy (anonymous):

2.56?

Vocaloid (vocaloid):

yes, so that means f(8) = 2.56 what is f(3)?

OpenStudy (anonymous):

0.08

Vocaloid (vocaloid):

right! f(8)=2.56 f(3) =0.08 so, average rate of change = [ f(8)-f(3) ] / (8-3) = ?

OpenStudy (anonymous):

0.496

Vocaloid (vocaloid):

good! that's our answer

OpenStudy (anonymous):

thx follow me at jherrera.29